# Logical Modalities within the European AI Act: An Analysis

### v1

@@ Abstract.
ccs: Computing methodologies Knowledge representation and reasoningccs: Applied computing Lawccs: Theory of computation Higher order logicccs: Theory of computation Automated reasoning
@@ 2.1. Obligations, Prohibitions, and Permissions
The AI Act is a European piece of legislation intended to regulate the use of AI systems. In it Obligations are expressed with the word ”shall”, as in this sentence from Article 8:
@@ 2.2. Contrary-to-Duty-Obligations
    If a system complies with the requirements, the provider must not take any corrective actions.
@@ 2.3. Agency and Agentive Obligations
is such an agentive obligatio because it expresses what the provider ought to do. Agentive obligations are different from general obligations, and must be distinguishable.
@@ 2.5. Temporality and Temporal Notions
This obligation for importers is tied to a specific point in time; it must be fulfilled before the system is placed on the market.
@@ 2.7. Fuzziness
2.8. Interactions and Combinations of Different Modalities
@@ 3. Which Logics?
To represent agentive obligations, a modal logic of agency is needed. While there are many ideas for formulating such a logic (Anderson, 1962; Fitch, 1963; Von Wright, 1963, 1981; Kanger, 1972; Brown, 1988), the most prominent theory of agency is a branch called Seeing-To-It-That (STIT) logic that originated in the works of Belnap (Belnap, 1991) and Belnap and Perloff (Belnap and Perloff, 1988). STIT theory introduces a STIT operator of the form
$stit_{a}F$
which expresses that an agent a sees to it that F holds. Given that agency appears in the AI Act mainly via agentive obligations, a suitable agency logic must be able to formulate ought-to-do statements for specific agents. Several authors have researched this issue and presented their approaches, which could be apt to help represent the AI Act (van Berkel and Lyon, 2019; Horty and Belnap, 1995; Horty, 2001; Murakami, 1998; Xu, 2015).
Moving to the representation of agent beliefs, another field in logic called epistemic logic becomes interesting. Epistemic logic allows for the exploration of different ideas of knowledge and beliefs of agents and their logical relations (Rendsvig et al., 2024). Popular epistemic logics include modal logics S4 and S5 based on possible world semantics, and providing modal operators representing what agents know and believe.
@@
Finally, a logic that can do justice to vague and fuzzy statements is called fuzzy logic. Fuzzy logic assigns a degree of truth to a proposition that is expressed on a scale within the real unit interval [0, 1], with 0 equal to total falsehood and one equal to total truth (Cintula et al., 2023). From the point of view of LogiKEy, however, the fuzzy logic variants of Hájek (Hájek, 1998) are particularly interesting, since they can be seen as generalized multi-value logics and as such are more amenable to formalization; in fact, (Hájek, 1998) provides a good blue-print for doing so in LogiKEy extending ideas from (Steen and Benzmüller, 2016).
@@ 4. Initial Experiments
In this section we give a brief summary of the findings and results from the experiments on embedding some of the logics using the LogiKEy approach and using these logics to represent parts of the AI Act; for details see (Author, 2024). Experiments have so far focused on only a selection of the identified modalities, namely obligations, prohibitions, permissions, CTDs, agency, and agentive obligations.
@@
To represent obligations (as well as permissions and prohibitions), SDL was identified as a suitable logic in Section 3. A trustworthy embedding of SDL in Isabelle/HOL using the LogiKEy approach already exists (Benzmüller and Parent, 2018) and has been used to represent parts from the AI Act, particularly Article 5 on prohibited AI systems. This was successful: The paragraphs could be adequately represented, and both Sledgehammer (an automated theorem proving tool integrated with Isabelle/HOL) and Nitpick (a model finder integrated with Isabelle/HOL) could reason with them correctly (Author, 2024). Sledgehammer was used to automatically check the validity of example judgements, and Nitpick to find countermodels for invalid judgements and also to prove the consistency of whole contexts.
@@
Like SDL, DDL already has a faithful embedding in Isabelle/HOL (Benzmüller et al., 2022). Using this embedding it was possible to represent selected CTDs from the AI Act without problems. The formalization of the selected examples from the AI Act was straightforward, and Nitpick and Sledgehammer could work with them successfully (Author, 2024). Since DDL can express everything SDL can and more (Gabbay et al., 2013), the DDL embedding is suitable for representing obligations, prohibitions, permissions, and CTDs. Encodings of Article 5 and selected CTDs from the AI Act in DDL can be found in the Appendix. [^1] Note: The appendix will be available in an arXiv-preprint version of this paper. All Isabelle files can be downloaded here: https://we.tl/t-veNrAlAJF4. Further work includes similar experiments adapting an existing embedding of Åqvist’s system E in Isabelle/HOL (Parent and Benzmüller, 2024).
@@
Additionally, logics for the representation of agency and agentive obligations were explored. In particular, an embedding into HOL of Temporal Deontic STIT Logic (van Berkel and Lyon, 2019), a logic appropriate for representing both actions of agents and agentive obligations, was created. While most axioms postulated in (van Berkel and Lyon, 2019) were provable via Sledgehammer and none were refuted, Nitpick could not find a model confirming the consistency of the embedding. This indicates that TDS logic has (presumably) only infinite models, making it unsuitable for model finding/checking in the context of any formalization of AI Act (Author, 2024) with existing tools.
@@
As an alternative to TDS, the authors explored the possibility of extending DDL with additional operators for the representation of agency and agentive obligations. Two different variants were introduced: One representing agent types as constants, with separate accessibility relations and obligation operators for each agent, and one representing agent types as function types, with a general accessibility relation and a general obligation operator taking an agent as an input parameter. Both variants introduced a STIT-like operator as a constant and equipped it with minimal axiomatization, namely to ensure that anything an agent sees to actually holds (Author, 2024).
@@
While the first variant is more restrictive, with agentive obligations always applying to a generic agent type without the possibility of specifying further constraints or relations, the second variant is more flexible, allowing groupings based on characteristics of agent types or relations and formulating obligations for specific subsets of agents. Unfortunately, the second variant performs worse in other aspects: It struggles with the representation of CTDs, and Nitpick can not find a model for cardinalities of i (domain of possible worlds) bigger than 1. For the first variant, a model is found up to a cardinality of i=2, and CTDs can be represented without problems. For both variants, all but one of the prominent lemmas for DDL as studied by Benzmüller et al. could be proven (Author, 2024). Whether either of the variants can be adapted or simplified to perform better remains an open question. The embedding of TDS and the two extensions of DDL can be found in the Appendix.
@@ 5. Conclusion
While the presented project is not yet completed, the reported material provides a starting point and references for future research. In the LogiKEy context, the challenge is to create embeddings of the suggested logics, choose and provide suitable logic combinations, and experiment with increasingly larger example parts of the AI Act.
@@
Acknowledgments:
Thanks to Geoff Sutcliffe for helpful comments and feedback on this document.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025conference: International Conference on Artificial Intelligence and Law; June 16–20, 2025; Chicago, ILccs: Computing methodologies Knowledge representation and reasoningccs: Applied computing Lawccs: Theory of computation Higher order logicccs: Theory of computation Automated reasoning
@@ 2.1. Obligations, Prohibitions, and Permissions
The AI Act is a European piece of legislation intended to regulate the use of AI systems. In it Obligations are expressed with the word ”shall”,  [^1] Within legal acts such as the AI Act, the world ”shall” can also have performative or constitutive functions instead of deontic ones. However, the primary function of ”shall” in the AI Act, is the introduction of deontic obligations. as in this sentence from Article 8:
@@ 2.2. Contrary-to-Duty-Obligations
    If a system complies with the requirements, the provider must not take any corrective actions.  [^2] It has been brought to our attention that an interpretation as ”need not” might be more appropriate. We are thankful for this remark and aim to uncover the differences between these interpretations in the future.
@@ 2.3. Agency and Agentive Obligations
is such an agentive obligation because it expresses what the provider ought to do. Agentive obligations are different from general obligations, and must be distinguishable.
@@ 2.5. Temporality and Temporal Notions
This obligation for importers is tied to a specific period and point in time; it must be fulfilled before the system is placed on the market.
@@ 2.7. Fuzziness
2.8. Combinations of Modalities
@@ 3. Which Logics?
To represent agentive obligations, a modal logic of agency is needed. While there are many ideas for formulating such a logic (Anderson, 1962; Fitch, 1963; Von Wright, 1963, 1981; Kanger, 1972; Brown, 1988), the most prominent theory of agency is a branch called Seeing-To-It-That (STIT) logic that originated in the works of Belnap (Belnap, 1991) and Belnap and Perloff (Belnap and Perloff, 1988). STIT theory introduces a STIT operator of the form $stit_{a}\ F$, which expresses that an agent a sees to it that F holds. Given that agency appears in the AI Act mainly via agentive obligations, a suitable agency logic must be able to formulate ought-to-do statements for specific agents. Several authors have researched this issue and presented their approaches, which could be apt to help represent the AI Act (van Berkel and Lyon, 2019; Horty and Belnap, 1995; Horty, 2001; Murakami, 1998; Xu, 2015).
Moving to the representation of agent beliefs, the field called epistemic logic becomes interesting. Epistemic logic allows for the exploration of different ideas of knowledge and beliefs of agents and their logical relations (Rendsvig et al., 2024). Popular epistemic logics include modal logics S4 and S5 based on possible world semantics, and providing modal operators representing what agents know and believe.
@@
Finally, a logic that can do justice to vague and fuzzy statements is called fuzzy logic. Fuzzy logic assigns a degree of truth to a proposition that is expressed on a scale within the real unit interval [0, 1], with 0 equal to total falsehood and one equal to total truth (Cintula et al., 2023). From the point of view of LogiKEy, however, the fuzzy logic variants of Hájek (Hájek, 1998) are particularly interesting, since they can be seen as generalized multi-value logics and as such are more amenable to formalization; in fact, (Hájek, 1998) provides a good blue-print for doing so in LogiKEy extending the formalization idea presented in (Steen and Benzmüller, 2016).
@@ 4. Initial Experiments
In this section we give a brief summary of the findings and results from the experiments on embedding some of the logics using the LogiKEy approach and using these logics to represent parts of the AI Act; for details see (Lawniczak, 2024). Experiments have so far focused on only a selection of the identified modalities, namely obligations, prohibitions, permissions, CTDs, agency, and agentive obligations.
@@
To represent obligations (as well as permissions and prohibitions), SDL was identified as a suitable logic in Section 3. A trustworthy embedding of SDL in Isabelle/HOL using the LogiKEy approach already exists (Benzmüller and Parent, 2018) and has been used to represent parts from the AI Act, particularly Article 5 on prohibited AI systems. This was successful: The paragraphs could be adequately represented, and both Sledgehammer (an automated theorem proving tool integrated with Isabelle/HOL) and Nitpick (Blanchette and Nipkow, 2010) (a model finder integrated with Isabelle/HOL) could reason with them correctly (Lawniczak, 2024). Sledgehammer was used to automatically check the validity of example judgements, and Nitpick to find countermodels for invalid judgements and also to prove the consistency of whole contexts.
@@
Like SDL, DDL already has a faithful embedding in Isabelle/HOL (Benzmüller et al., 2022). Using this embedding it was possible to represent selected CTDs from the AI Act without problems. The formalization of the selected examples from the AI Act was straightforward, and Nitpick and Sledgehammer could work with them successfully (Lawniczak, 2024). Since DDL can express everything SDL can and more (Gabbay et al., 2013), the DDL embedding is suitable for representing obligations, prohibitions, permissions, and CTDs. Encodings of Article 5 and selected CTDs from the AI Act in DDL can be found in the appendix. [^3] The Isabelle/HOL source files can also be downloaded here: http://logikey.org/tree/master/2025-ICAIL-Data. Further work includes similar experiments adapting an existing embedding of Åqvist’s system E in Isabelle/HOL (Parent and Benzmüller, 2024).
@@
Additionally, logics for the representation of agency and agentive obligations were explored. In particular, an embedding into HOL of Temporal Deontic STIT Logic (van Berkel and Lyon, 2019), a logic appropriate for representing both actions of agents and agentive obligations, was created. While most axioms postulated in (van Berkel and Lyon, 2019) were provable via Sledgehammer and none were refuted, Nitpick could not find a model confirming the consistency of the embedding. This indicates that TDS logic has (presumably) only infinite models, making it unsuitable for model finding/checking in the context of any formalization of AI Act (Lawniczak, 2024) with existing tools.
@@
As an alternative to TDS, the authors explored the possibility of extending DDL with additional operators for the representation of agency and agentive obligations. Two different variants were introduced: One representing agent types as constants, with separate accessibility relations and obligation operators for each agent, and one representing agent types as function types, with a general accessibility relation and a general obligation operator taking an agent as an input parameter. Both variants introduced a STIT-like operator as a constant and equipped it with minimal axiomatization, namely to ensure that anything an agent sees to actually holds (Lawniczak, 2024).
@@
While the first variant is more restrictive, with agentive obligations always applying to a generic agent type without the possibility of specifying further constraints or relations, the second variant is more flexible, allowing groupings based on characteristics of agent types or relations and formulating obligations for specific subsets of agents. Unfortunately, the second variant performs worse in other aspects: It struggles with the representation of CTDs, and Nitpick can not find a model for cardinalities of i (domain of possible worlds) bigger than 1. For the first variant, a model is found up to a cardinality of i=2, and CTDs can be represented without problems. For both variants, all but one of the prominent lemmas for DDL as studied by Benzmüller et al. could be proven (Lawniczak, 2024). Whether either of the variants can be adapted or simplified to perform better remains an open question. The embedding of TDS and the two extensions of DDL can be found in the appendix.
@@ 5. Conclusion
While the presented project is not yet completed, the reported material provides a starting point and references for future research. In the LogiKEy context, the challenge is to create embeddings of the suggested logics, choose and provide suitable logic combinations, and experiment with increasingly larger parts of the AI Act.
@@
Acknowledgements
We are grateful to the anonymous reviewers for their helpful feedback and comments.